UARS Weekly Status Report

01 September 2000




********************************************************************

                UARS WEEKLY STATUS REPORT - 01 SEP 2000

********************************************************************

***********
SOLSTICE
***********

SOLSTICE is in good health, collecting solar data on all available orbital
tracking periods and stellar calibration observations whenever possible.
Our solar data is not affected by the larger than normal pointing errors
(2-4 arcminutes across a track) since they are within our calibrations.
Additionally, we continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
Level2 processing is halted due to bugs in the OA_SSPP_ATT service. CDHF
and CSC personnel are working on the problem.

***********
SUSIM
***********

Nothing received.

***********
ACRIM
***********

The ACRIM II experiment continues to gather data normally. There has been 
no impact on data acquisition by the loss of star tracker information. Data 
processing for current data is up to date and re-processing of all ACRIM II 
data is proceeding on schedule. A comprehensive review of sensor 
degradation/calibration for the duration of the mission is under way as 
part of the re-processing activity.

***********
PEM
***********

Nothing received.

***********
HALOE 
***********

Nothing received.

***********
MLS
***********

Nothing received.

***********
HRDI
***********

Nothing received.

***********
WINDII
***********

Nothing received

***********
MPG
***********

Processing of instrument Daily Activity Plans continues with no
problems. We are continuing to coordinate with the Flight
Operations Team and to assist the Project Scientist in the
scheduling of instrument operations.

The Instrument Operations Schedule for the yaw period
17 September to 22 October 2000 was distributed.

***********
CDHF
***********

The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
   (C) = Complete, (F)= Failed  (L) = Days remaining to reprocess.
     -  No reprocessing requested this reporting period.
                   
Data created in the month of July, 2000 will be previewed and loaded
into the Daac Transfer Tables September, 2000.
                        
UARS CDHF REHOST:
   -  Currently investigating differences noted in the SUSIM output file
      sizes for jobs run on the current (PROD) vs. rehost (TEST) systems.       
   -  Continue to do differences and compare LEVEL0 ingested on the
      current and rehost systems. Results all are good.
   -  Continue to monitor the current and rehost systems for
      CORRELATIVE (UKMO) and Auxiliary (SOL) data that are ingested daily.
    ISSUES:  
    1. Positive response from all PI End Users prior to transition.
    2. PEM not Alpha compliant. PEM forwarded (or will be forwarded)
         version we believe can process on Alpha.
    3. Migration of DAAC and FORMS tables.   
                  
***********
FOT
***********

                                 REPORTING PERIOD

This report covers 19 August 2000 (Orbit 48857, GMT Day 232) through 25 August 
2000 (Orbit 48961, GMT Day 238).

                               SPACECRAFT OPERATION

The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is 
performing nominally under two-battery operations.  The instrument operational 
changes were:


                          INSTRUMENT OPERATIONAL CHANGES


            Instrument     Time        Orbit       Comment
            HALOE        232/0033      48857       Power ON

The beta angle ranged between the angles of 15.5 degrees to 39.5 degrees for 
this report period.  The beta angle is now increasing toward a maximum angle of 
+49 degrees on 31 August 2000 (DOY 244).

Spacecraft battery 1, 2, and 3 performance monitoring continues.  Battery 1 
remains OFF the charge relay.  Battery 2 and Battery 3 are currently 
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.  Battery 
temperatures are stable with a temperature delta between Battery 2 and Battery 
3 of 0.6 to 0.8 degrees C.  Additional battery performance data for each 
battery on  18 August 2000 and 25 August 2000 are provided in Appendix D.

The clock error ranged from -10.3 to 10.1 msec during this report period.  
Clock rate adjustments are listed in Appendix C.

All instrument and subsystem engineers were advised on the initial excursion of 
ALERTS or Out-Of-Limits conditions detected by the FOT for this report.  A 
summary is listed below and the details of ALERT and Out-Of-Limits occurrences 
are listed in Appendix B.

     OUT-OF-LIMITS  =  00
     ALERTS         =  00

The following real-time command sheets were executed for the listed operational 
element(s) during this report period (see Appendix A).

     FOT            =  35
     HALOE          =  02

UARS Anomaly Reports initiated during this report period are listed in Appendix 
E and summarized below.
 
     FOT            =  01

Three (3) Anomaly Reports remained OPEN at the end of this report period (see 
Appendix E).

TRANSPONDER FREQUENCIES:
 Transponder A:  00/153  -  2287.496961
 Transponder B:  00/213  -  2287.499288


                          GROUND SYSTEM OPERATION

One (1) operational support problem occurred this period.  Anomaly Reports are 
included in Appendix E.

The data loss calculations supplied by Data Capture Facility (DCF) from mission 
start through UTC Day 00/237 show the observatory data loss to be 298 hours, 11 
minutes, 23.369 seconds (an increase of 18 minutes 44.184 seconds since the 
last report period).  This is a 0.40 percent data loss which equals a 99.60 
percent data capture for the mission.

The increase of 18 minutes 44.184 seconds occurred over four (4) days:

 00 minutes 27.328 seconds on Day 231 due to a service interruption,
 00 minutes 12.360 seconds on Day 232 due to a data hit,
 01 minutes 15.792 seconds on Day 234 due to a late acquisition,
 16 minutes 48.704 seconds on Day 236 due to data hits.


                         OTHER SIGNIFICANT EVENTS

                         UARS SATELLITE OPERATIONS

The FOT continues to operate FHST-1 for approximately one orbit each day.  By 
cycling the tracker on and off for such a short duration the degradation seems 
to have been reduced.  Scale factor calibrations are being provided on a 
regular basis from FDF and are being uplinked when received.  It appears that 
the FOT has been able to stabilize the performance of FHST-1 for now and 
continues to work on establishing another viable attitude control solution 
should the tracker fail.

                                MEETINGS

The FOT attended a meeting to discuss FHST-1 on 21 and 23 August 2000.

                        FUTURE SIGNIFICANT EVENTS

The next UARS Yaw slew is scheduled for September 17, 2000 (DOY 261).  This 
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere 
viewing).

Attachments: Appendix A,  Real-time Commands Sheets
             Appendix B,  Out-Of-Limits & ALERT Occurrences
             Appendix C,  Clock & Solar Array Rate Corrections
             Appendix D,  UARS Battery Performance Data
             Appendix E,  UARS Anomaly Reports

                                 APPENDIX A
                           REAL-TIME COMMAND SHEETS
                           Week ending 25 Aug 2000

ORBIT     AOS          NO.        SUBSYS     REASON
-----     ---          ---        ------     -------
48857     232/0033     00-425     HALOE      Power ON (Part-1)
48859     232/0406     00-423     NBTR       PLAYBACK
48860     232/0628     00-423     NBTR       NOOPS D-232
48860     232/0628     00-426     HALOE      Power ON (Part-2)
48863     232/1133     00-423     NBTR       NOOPS D-232
48866     232/1604     00-422     FHST-1     Power ON
48867     232/1749     00-422     FHST-1     Power OFF
48868     232/1829     00-427     NBTR       NOOPS D-233
48874     233/0409     00-427     NBTR       PLAYBACK
48875     233/0553     00-427     NBTR       NOOPS D-233
48881     233/1607     00-422     FHST-1     Power ON
48882     233/1751     00-422     FHST-1     Power OFF
48883     233/1831     00-428     NBTR       NOOPS D-234
48889     234/0412     00-428     NBTR       PLAYBACK
48890     234/0555     00-428     NBTR       NOOPS D-234
48896     234/1610     00-422     FHST-1     Power ON
48897     234/1753     00-422     FHST-1     Power OFF
48905     235/0557     00-429     NBTR       NOOPS D-235
48911     235/1506     00-422     FHST-1     Power ON
48913     235/1937     00-422     FHST-1     Power OFF
48919     236/0454     00-430     NBTR       DOY 236 NBTR NOOPS
48926     236/1615     00-422     FHST-1     Power ON
48927     236/1757     00-431     NBTR       NOOPS D-237
48927     236/1757     00-422     FHST-1     Power OFF
48929     236/2120     00-432     FHST-1     CALIBRATION
48932     237/0046     00-433     OBC        Clk Adj fr 0.00 to 20.6 msec/day
48934     237/0501     00-431     NBTR       NOOPS D-237
48941     237/1617     00-422     FHST-1     Power ON
48942     237/1758     00-422     FHST-1     Power OFF
48943     237/1942     00-434     NBTR       NOOPS D-434
48949     238/0458     00-434     NBTR       DOY 238 NBTR NOOPS
48953     238/1145     00-435     OBC        Clk Adj fr 20.6 to 0.00 msec/day
48956     238/1619     00-422     FHST1      Power ON
48956     238/1619     00-434     NBTR       NOOPS D238
48957     238/1800     00-436     NBTR       NOOPS D239
48957     238/1800     00-422     FHST1       Power OFF
48958     238/1844     00-437     FHST1       SCALE FACTOR UPDATE
_______________________________________________________________________________

                                APPENDIX B
                     OUT-OF-LIMITS & ALERTS OCCURRENCES
                         Week ending 25 Aug 2000

The following Out-Of-Limits occurrences were detected:

MNEMONIC      STATE          DESCRIPTION           ORBITS
---------     -----          ------------          ------
              -None this report period


The following ALERT occurrences were detected:

MNEMONIC      DESCRIPTION                          ORBITS
---------     ------------                         ------
             -None this report period

_______________________________________________________________________________

                                APPENDIX C
                   CLOCK & SOLAR ARRAY RATE CORRECTIONS
                          Week ending 25 Aug 2000

                        CLOCK DRIFT RATE ADJUSTMENTS

ORBIT       TIME                      CLOCK RATE CHANGE
-----       ----                      -----------------
48932       237/0050 (24 August)      from 0.00 to 20.6 msec/day
48953       238/1148 (25 August)      from 20.6 to 0.00 msec/day
___________________________________________________________________________

                                APPENDIX D
                       UARS BATTERY PERFORMANCE DATA
                           Week ending 25 Aug 2000


     18 Aug 2000 - GMT Day 231 Beta = 15.5 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)     CURR (amp)    EON V     AVE C/D   AVE
BAT  MAX/MIN        MAX/MIN      MAX/MIN       MIN       RATIO     DOD(%)
---  -----------    ---------    ----------    ------    -------  ------
1    +459.2/+168.0  -3.05/-4.62  +0.0/+0.0     15.2     0.000     0.0
2    +33.6/-44.8    +5.79/+3.87  +29.6/-14.4   24.3     1.040     19.6
3    +22.4/-33.6    +4.83/+2.60  +28.8/-15.2   24.3     1.037     19.2


     25 Aug 2000 - GMT Day 238 Beta = 39.5 deg, SA Pos = 269 deg

     DIFF V (mV)    TEMP (C)     CURR (amp)    EON V    AVE C/D   AVE
BAT  MAX/MIN        MAX/MIN      MAX/MIN       MIN      RATIO     DOD(%)
---  -----------    ---------    -----------   ------   -------   ------
1    +728.0/+728.0  -3.05/-3.99  +0.0/+0.0    .14.4     0.000     0.0
2    +39.2/-67.2    +5.79/+4.51  +28.0/-15.2   24.5     1.030     21.4
3    +22.4/-50.4    +5.15/+3.87  +27.6/-15.6   24.5     1.030     21.3


     V/T LEVEL CHANGES
ORB#      TIME              LVL FR-TO
-----     ----------        ----------
          no changes        V/T 5
_______________________________________________________________________________
_
                                  APPENDIX E
                           UARS ANOMALY REPORTS
                          Week ending 25 Aug 2000

                       NEW ANOMALY REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME       TTR      PROBLEM/DATA LOSS
------     ------     -----     ----       ---      -----------------
00-037     FOT         48898    234/1833   14170    Generic Late Acq.

                   ANOMALY CLOSURE INFORMATION RECEIVED
AR NO.     SUBSYS     ORBIT     TIME       TTR      PROBLEM/DATA LOSS
------     ------     -----     ----       ---      -----------------
            -None this report period

                 ANOMALY INVESTIGATION REPORTS GENERATED
AR NO.     SUBSYS     ORBIT     TIME       TTR      PROBLEM/DATA LOSS
------     ------     -----     ----       ---      -----------------
             -None this report period

                       OPEN ANOMALY REPORTS
AR NO.     SUBSYS     ORBIT     TIME       TTR      PROBLEM/DATA LOSS
------     ------     -----     ----       ---      -----------------
00-040     FOT        48854     231/2111   22831    Data Dropout
00-037     FOT        see list  Aug 00     14170    Generic Late Acq.
00-032     FHST       48211     188/2242   N/A      FHST 2 Anomaly

	 	SEVNTFW00/34